Find the value of r in (4,r), (r,2) so that the slope of the line containing them is -5/3.
Answer by nerdybill(7384)

Slope of between any two points is given by:
m = (y2-y1)/(x2-x1)
.
In your case, you were given:
(x1,y1) = (4,r)
(x2,y2) = (r,2)
m = -5/3
.
Plugging the above into:
m = (y2-y1)/(x2-x1)
we get
-5/3 = (2-r)/(r-4)
Now, we solve for r:
.
-5/3 = (2-r)/(r-4)
-5 = 3(2-r)/(r-4)
-5(r-4) = 3(2-r)
-5r+20 = 6-3r
20 = 6+2r
14 = 2r
7 = r